OpenStudy (anonymous):

In an election, 11 of 20 voters preferred Ms. Williamson to Ms. Stallings. A total of 3300 people voted in the election. How many voters preferred Ms. Williamson to Ms. Stallings? A.1815 B.3291 C.4185 D.6000 Please help me

OpenStudy (anonymous):

@Teddyiswatshecallsme @texaschic@autumn421-eva @pooja195 @LeeEtchison please help!

OpenStudy (anonymous):

this is confusing

OpenStudy (anonymous):

as you can tell i am not the best at math

OpenStudy (anonymous):

@autumn421-eva

OpenStudy (anonymous):

this is still proportion applications

OpenStudy (anonymous):

11/20= x/3300

OpenStudy (anonymous):

oh, so do that then i will get the answer??

OpenStudy (anonymous):

hang on a sec

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i will tag you back when i am done

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{11\ 3300}{20}=1815 \]

OpenStudy (anonymous):

@autumn421-eva

OpenStudy (anonymous):

i think it is 1815, and not because robtobey said so

OpenStudy (anonymous):

i really did it

OpenStudy (anonymous):

that is correct

OpenStudy (anonymous):

thx